OpenStudy (anonymous):

Find the derivative of f(x) = -(9/x) at x = -8

OpenStudy (anonymous):

so u can rewrite that as f(x)=-(9*x^(-1)) and when u find derivative of that just replace x for -8

OpenStudy (anonymous):

but my options are A. 8/9 B. 9/8 C. 64/9 D. 9/64

OpenStudy (anonymous):

do you know how to find derivative of -(9*x^(-1))?

OpenStudy (anonymous):

Don't I plug in (x+h) where x is?

OpenStudy (anonymous):

ok so we need derivative of -(9/x) 1. we can plug out of the the -9 since its a constant and look for derivative of only 1/x can u find derivative od 1/x?

OpenStudy (anonymous):

ummm would it just be 1/-8?

OpenStudy (anonymous):

ok so we can write 1/x as x^(-1) the derivative of that we can fine by the formula x^a=ax^(a-1) , so we would have -1*x^(-2) which is same as 1/x^2 and multiplied by the constant we put away before -9 it would be 9/x^2 now just replace for x=-8 in 9/x^2 and u have the answer

OpenStudy (anonymous):

ohhhh, jeez why can't teachers just explain it that way. Its so much easier. Thanks so much man!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

can you explain one more for me? well help me figure it out

OpenStudy (anonymous):

yeah sure go ahead

OpenStudy (anonymous):

Find the derivative of f(x) = -10x2 + 4x at x = 11.

OpenStudy (anonymous):

my teacher told me the with the x I plug in (x+h)

OpenStudy (anonymous):

is that true?

OpenStudy (anonymous):

I kind of need a step by step with this one

OpenStudy (anonymous):

um i dont really know the way with the plug in (x+h) if he told you probably its correct but cant remember really , if you want i can solve but not in that way

OpenStudy (anonymous):

alright that would be fine, let me see what you get

OpenStudy (anonymous):

its quite similar to the above using same formula if f(x)=x^a where a is any number then f'(x)=a*x^(a-1), so basicaly u need to use that twice for each term means for -10x^2 and +4x

OpenStudy (anonymous):

so I plug in 11 at each x?

OpenStudy (anonymous):

so lets go first for -10*x^2 same we can put the -10 aside since its constant and look for derivative of x^2

OpenStudy (anonymous):

using the formula a*x^(a-1) for a=2 we would get 2*x^1 but we dont write the 1 so its only 2x

OpenStudy (anonymous):

and since we put the -10 aside now we need to multiply it by -10 which leae us -20x for the first term

OpenStudy (anonymous):

okay so now we have -20x + 4x?

OpenStudy (anonymous):

no now you need to fine derivative of the second term means derivative of 4x

OpenStudy (anonymous):

you do it same way but now for a=1 (because x is of degree 1)

OpenStudy (anonymous):

so do you get 1 or do you get x^0?

OpenStudy (anonymous):

so yeah for the x you would get x^0 which equals 1(since any number in power of 0 is 1) but dont forget to multiply it with the constant 4

OpenStudy (anonymous):

so now I have -20 + 4?

OpenStudy (anonymous):

for the first one we had -20x , so you get -20x+4 , and now you only left to substitude for the value of x you need which was 11 i belive

OpenStudy (anonymous):

oh so I just plug in the x now so i get -220 + 44?

OpenStudy (anonymous):

176 is my answer?

OpenStudy (anonymous):

i dont know where u get the 44 from , because u dont have 4x you have only 4 so you dont multiply that by 11 so would leave you -220+4

OpenStudy (anonymous):

oh yeahh its just 4... whoops! but thanks! that formula is really nice!

OpenStudy (anonymous):

no problem , ur welcomed ^^